rxcklesskaisher:

Lisa has to cut her grandma's grass this weekend and wants to know exactly how much area she will be cutting. Calculate the area of the polygon. Be sure to show all your work and explain your answer. (typing answer)

rxcklesskaisher:

1 attachment
AZ:

To find the area of the polygon, let's split it into two different trapezoids like this: I also went ahead and wrote down the side lengths from the information they gave you

1 attachment
AZ:

\( \sf Area ~of~a~ trapezoid = \left(\dfrac{base_1 + base_2}{2}\right) \times height\)

AZ:

What is the area of the first trapezoid?  Base 1 is 26 Base 2 is 6 and the height is 20

AZ:

What is the area of the second trapezoid?  Base 1 is 16 Base 2 is 10 and the height is 6

AZ:

And then to find the total area of the polygon, just add up the two numbers you got because the area of the polygon is the area of both the trapezoids added together